# Coupling of magnetic and lattice collective excitations in the 2D van der Waals antiferromagnet FePS3

### v1

@@ Abstract
We report a comprehensive magneto-optical study of FePS3 single crystals using temperature-dependent polarized transmission spectroscopy and Faraday rotation measurements. Our data reveal strong in-plane anisotropy of both phonon and magnon excitations. These anisotropies emerge distinctly below the Néel temperature ($T_{N}\approx 118\leavevmode\nobreak\ K$). A prominent magnon mode at 122  $cm^{-1}$ exhibits characteristic redshifting with temperature and field-induced splitting, confirming its magnetic origin. The optical conductivity, extracted via Drude–Lorentz modeling, displays polarization-sensitive spectral changes across the magnetic transition. The observation of Raman-active modes in the infrared spectra and the detection of multiple phonons exhibiting Faraday rotation provide compelling evidence of spin–phonon coupling and inversion symmetry breaking in the magnetic state. These findings underscore the intricate interplay between magnetic and lattice degrees of freedom in FePS3 and establish its relevance for exploring spin–lattice interactions in two-dimensional antiferromagnets.
@@ I Introduction
Two-dimensional magnets are also an unparalleled platform for testing the fundamental physics of spin models. They permit direct access to Ising [5], XY [6, 7], Kitaev, and Heisenberg-type physics within a single atomic layer. This allows one to observe, manipulate, and exploit fluctuations of the magnetic order, and to probe how external stimuli modify the ground state [8, 9, 10, 11]. Robust (anti)ferromagnetic order at the 2D limit can emerge in the presence of magnetic anisotropy, interlayer coupling, or engineered perturbations. In practice, the resulting order and its excitations are exquisitely sensitive to the spin dimensionality, crystal-field anisotropy, and external tuning parameters such as strain, doping, and defects.
@@
Magnetic van der Waals materials, including metal phosphorus trichalcogenides (MPX3) [M=Fe, Co, Ni, Mn; X=S, Se], represent a broad class of compounds with diverse spin dimensionality models, exhibiting both ferromagnetic and antiferromagnetic states. For example, FePS3 ($T_{\mathrm{N}}\approx 120$ K), MnPS3 ($T_{\mathrm{N}}\approx 78$ K), and NiPS3 ($T_{\mathrm{N}}\approx 155$ K), all with monoclinic crystal structures (space group C2/m), display distinct magnetic behaviors [12, 13]. FePS3 adheres to an Ising model with a zigzag antiferromagnetic configuration, featuring in-plane ferromagnetic chains that are antiparallel along one of the in-plane crystal axes, while NiPS3 and MnPS3 are best described by a XY and Heisenberg model, respectively [12, 13]. MnPS3 shows a Néel-type antiferromagnetic order, where neighboring atoms have antiparallel spins, in contrast to the zig-zag antiferromagnetic structure in FePS3. Notably, FePS3 maintains its magnetic order at the monolayer limit [14, 15] and undergoes an insulator-to-metal transition under quasi-hydrostatic pressure [16]. Inelastic neutron scattering measurements of FePS3 have identified two magnon excitations with finite gaps of 15 meV and 40 meV at $q=0$ [17]. Moreover, the comparable magnitudes of spin-orbit coupling and crystal field energies in FePS3 favor the presence of low-energy magnetic excitations [18]. Yet, a comprehensive understanding of how these low-energy magnetic excitations interact with phonons, what their circular conductivity spectral weight is, and how they affect crystal symmetries in FePS3 is still lacking.
@@
The magnetic properties of these vdW materials are strongly coupled to the charge and lattice degrees of freedom and can be studied effectively using far-infrared spectroscopy. The additional advantage of optical spectroscopy is that it is also susceptible to the monolayer limit. Therefore, the magnetic state can be easily studied and detected using optical probes, such as Kerr/Faraday angle rotation, compared to other magnetometry techniques [19, 20]. The advantage of infrared spectroscopy is the direct access to the spectral weight of the different collective modes in the magnetic phase and, as a result, to the investigation of their coupling and relations to the underlying physics of the magnetic state.
@@
In this article, we present a comprehensive infrared spectroscopy investigation of the 2D antiferromagnetic semiconductor FePS3. Using polarized infrared transmission spectroscopy and Faraday rotation measurements, we aim to elucidate the temperature and magnetic field dependence of both phonon and magnon excitations. Our primary objectives are to probe anisotropic optical responses, detect potential inversion symmetry breaking, and clarify the interplay between spin and lattice degrees of freedom in FePS3. To address these objectives, we synthesized high-quality FePS3 crystals and performed detailed magneto-optical measurements as described below.
@@ II.2 Optical spectroscopy
To measure the magneto-optical coefficients, we have used a Bruker Vertex 70v Fourier Transform Infra-Red (FTIR) spectrometer coupled to a Cryogenic magneto-optical cryostat, allowing transmission and reflection measurements as a function of temperature, magnetic field, and incoming/outgoing polarization. We have used Hg arc lamp and Globar (SiC) sources for the far-infrared (FIR) and mid-infrared (MIR) frequency ranges, respectively. Residual magnetic fields affect the arc lamp, limiting reliable measurements below 40 $cm^{-1}$. Therefore, we primarily present results obtained with the Globar source. To detect the transmitted signal intensity in our measurements, we have used several liquid-helium-cooled Si bolometers from Infra-Red Labs, covering the range from about 10 cm-1 to about 4000 cm-1. The magnetic field of the cryostat can be tuned from -7 T to +7 T and was aligned parallel (perpendicular) to the light propagation vector (sample surface normal) in the Faraday configuration. The temperature range of the cryostat sample holder is from 4 K to about 300 K, while the sample is mounted with silver paste to the sample holder and situated in the vacuum chamber as the superconducting magnet during the measurement. The cryostat is positioned on an external linear translation stage, allowing the measurement of a sample in transmission (reflection) configuration and a bare hole (mirror) reference at any desired temperature, magnetic field, and incoming/outgoing polarization.
@@
For the polarization-dependent measurements, we have installed gold or KRS polarizers for the FIR or MIR frequency ranges, respectively, along the incoming and outgoing light beam paths. The incoming polarizer was used for the temperature and magnetic field-dependent measurement for each principal polarization axis without an outgoing analyzer. In the Faraday rotation angle measurements, we fix the analyzer and rotate the incoming polarizer. Since FePS3 has a monoclinic crystal structure and shows an anisotropic optical response in the antiferromagnetic state, the typical fast protocol procedure [22], taking into account only four measurements, i.e. combination of external magnetic field ±$\pm B$ and $A=P=\pm 45^{\circ}$, where $P$ and $A$ are the incoming polarizer and outgoing analyzer angles, respectively, is subject to issues of numerical error. Therefore, we have applied the full protocol to measure the Faraday rotation angle for each of the two principal polarizations used in this work. In this protocol, the analyzer angle is fixed to one of the principal angles while measurement of the transmission intensity is being taken at intervals of $15^{\circ}$ for a full $360^{\circ}$ polarizer rotation. We have conducted the full protocol at a temperature of 5 K, i.e., deep in the antiferromagnetic state, and at fields of $B=\pm 7\leavevmode\nobreak\ T$. In addition, we have applied the fast protocol to investigate the temperature and magnetic field evolution of the Faraday angle rotation in FePS3.
@@ III.1 Infrared Transmission
In Figure 1, we show the transmission color map measured along the two principal axes. On one hand, the transmission spectrum at temperatures substantially higher than the Néel antiferromagnetic transition temperature, $T_{N}=118\leavevmode\nobreak\ K$, of  FePS3 shows no major difference between the two principal axes (See in App. VI). On the other hand, a clear change in the spectra can be seen below the phase transition. Previous attempts to fit the transmission data of the non-polarized spectrum and therefore to extract the optical conductivity have failed, indicating the importance of the polarization dependence of the transmission spectrum and the anisotropy of the monoclinic crystal of  FePS3 in the magnetic state.
Figure 1: Polarized transmission through a 26 $\mu$m FePS3 flake as a function of frequency and temperature for two principal polarization (upper panel: Polarization 1, and lower panel: Polarization 2). A pronounced spectral evolution occurs below $T_{N}=118$ K, including the emergence of new phonons and redshifting of the 122 cm-1 magnon mode.
@@
Table 1: Comparison of theoretical and experimental (Pol 1 and Pol 2) vibrational modes in FePS3. All values are in cm-1.
Symmetry Theoretical Pol 1 Pol 2
$\mathrm{B}_{u}$(1) 75.8
$\mathrm{B}_{u}$(2) 101.9 102.1
107.7
$\mathrm{B}_{u}$(3) 119.4
$\mathrm{A}_{u}$(1) 129.7 128.5 133.1
$\mathrm{A}_{u}$(2) 139.0 140.8 140.8
@@
$\mathrm{A}_{u}$(3) 153.8 153.1
160.3
$\mathrm{B}_{u}$(4) 161.7 163.5
$\mathrm{B}_{u}$(5) 167.6 165.2 165.9
$\mathrm{B}_{u}$(6) 174.5 175.7 174.5
$\mathrm{A}_{u}$(4) 178.0
$\mathrm{B}_{u}$(7) 179.3 184.7 184.8
$\mathrm{B}_{u}$(8) 202.5
$\mathrm{A}_{u}$(5) 202.7
$\mathrm{A}_{u}$(6) 207.5 207.4
$\mathrm{A}_{u}$(7) 221.6 215.4 215.1
$\mathrm{B}_{u}$(9) 230.3 228.0 227.7
$\mathrm{B}_{u}$(10) 231.9 232.2
$\mathrm{B}_{u}$(11) 240.1
$\mathrm{B}_{u}$(12) 254.6 257.4 260.1
$\mathrm{B}_{u}$(13) 350.2 375.6
$\mathrm{B}_{u}$(14) 415.1
$\mathrm{B}_{u}$(15) 531.4 531.9
$\mathrm{A}_{u}$(8) 538.9
$\mathrm{B}_{u}$(16) 541.7 546.1
$\mathrm{B}_{u}$(17) 550.3
$\mathrm{A}_{u}$(9) 556.7 556.3 555.0
Along with the drastic change of the overall transmission spectra below $T_{N}$, a new mode appears at about 122  $cm^{-1}$, which seems to harden as temperature is lowered. We will now focus on the spectral weight, temperature, magnetic field, and polarization-induced changes in the collective excitations spectrum, which are coupled to the magnetic order.

### v2

@@ Abstract
We combine polarized infrared magneto-transmission and Faraday angle rotation measurement to map the collective excitations of the van der Waals antiferromagnet FePS3. Below the Néel temperature ($T_{\mathrm{N}}\approx 118\leavevmode\nobreak\ K$), the phonon spectrum becomes strongly anisotropic, while a prominent excitation at $122$  $cm^{-1}$ ($15$ meV) hardens on cooling and splits linearly with magnetic field, identifying its magnetic origin. From absolute transmission and the Faraday angle, we reconstruct the circular optical conductivities, revealing a pronounced dichroism of the field-split excitations. The upper branch near $129$  $cm^{-1}$ shows an anomalously reduced dichroism consistent with spin–phonon hybridization with a nearby infrared phonon. Several phonons exhibit sizable Faraday rotation, and Raman-active modes appear in the infrared spectra; together with a broad mid-infrared band near $900$  $cm^{-1}$ emerging only below $T_{\mathrm{N}}$, these observations point to inversion-symmetry breaking in the magnetic state. A weak Faraday feature near $320$  $cm^{-1}$ ($\sim 40$ meV) further supports the presence of a higher-energy magnetic excitation. These findings underscore the intricate interplay between magnetic and lattice degrees of freedom in FePS3 and establish its relevance for exploring spin–lattice interactions in two-dimensional antiferromagnetic materials.
@@ I Introduction
Two-dimensional magnets are also an unparalleled platform for testing the fundamental physics of spin models. They permit direct access to Ising [5], XY [6, 7], Kitaev, and Heisenberg-type physics within a single atomic layer. This allows one to observe, manipulate, and exploit fluctuations of the magnetic order, and to probe how external stimuli such as temperature, field, strain, pressure modify the ground state [8, 9, 10, 11]. Robust (anti)ferromagnetic order at the 2D limit can emerge in the presence of magnetic anisotropy, interlayer coupling, or engineered perturbations. In practice, the resulting order and its excitations are exquisitely sensitive to the spin dimensionality, crystal-field anisotropy, and external tuning parameters such as strain, doping, and defects.
@@
Magnetic van der Waals materials, including metal phosphorus trichalcogenides (MPX3) [M=Fe, Co, Ni, Mn; X=S, Se], represent a broad class of compounds with diverse spin dimensionality models, exhibiting both ferromagnetic and antiferromagnetic states. For example, FePS3 ($T_{\mathrm{N}}\approx 120$ K), MnPS3 ($T_{\mathrm{N}}\approx 78$ K), and NiPS3 ($T_{\mathrm{N}}\approx 155$ K), all with monoclinic crystal structures (space group C2/m), display distinct magnetic behaviors [12, 13]. FePS3 adheres to an Ising model with a zigzag antiferromagnetic configuration, featuring in-plane ferromagnetic chains that are antiparallel along one of the in-plane crystal axes, while NiPS3 and MnPS3 are best described by a XY and Heisenberg model, respectively [12, 13]. MnPS3 shows a Néel-type antiferromagnetic order, where neighboring atoms have antiparallel spins, in contrast to the zigzag antiferromagnetic structure in FePS3. Notably, FePS3 maintains its magnetic order at the monolayer limit [14, 15] and undergoes an insulator-to-metal transition under quasi-hydrostatic pressure [16]. Inelastic neutron scattering measurements of FePS3 have identified two excitations with finite gaps of 15 meV and 40 meV at $q=0$ [17]. Moreover, the comparable magnitudes of spin-orbit coupling and crystal field energies in FePS3 favor the presence of low-energy magnetic excitations [18]. Yet, a comprehensive understanding of how these low-energy magnetic excitations interact with phonons, what their circular conductivity spectral weight is, and how they affect crystal symmetries in FePS3 is still lacking.
@@
The magnetic properties of these vdW materials are strongly coupled to the charge and lattice degrees of freedom and can be studied effectively using far-infrared spectroscopy. The advantage of optical spectroscopy is that it is also susceptible to the monolayer limit. Therefore, the magnetic state can be easily studied and detected using optical probes, such as Kerr/Faraday angle rotation, compared to other magnetometry techniques [19, 20]. The additional advantage of infrared spectroscopy is the direct access to the spectral weight of the different collective modes in the magnetic phase and, as a result, to the investigation of their coupling and relations to the underlying physics of the magnetic state.
@@
In this article, we present a comprehensive infrared spectroscopy investigation of the 2D antiferromagnetic semiconductor FePS3. Using polarized infrared transmission spectroscopy and Faraday rotation measurements, we aim to elucidate the temperature and magnetic field dependence of the collective modes’ spectra. Our primary objectives are to probe anisotropic optical responses, detect potential inversion symmetry breaking, and clarify the interplay between spin and lattice degrees of freedom in FePS3. To address these objectives, we synthesized high-quality FePS3 crystals and performed detailed magneto-optical measurements as described below.
@@ II.2 Infrared spectroscopy
To measure the magneto-optical coefficients, we have used a Bruker Vertex 70v Fourier Transform Infra-Red (FTIR) spectrometer coupled to a Cryogenic magneto-optical cryostat, allowing transmission measurements as a function of temperature, magnetic field, and incoming/outgoing polarization. We have used Hg arc lamp and Globar (SiC) sources for the far-infrared (FIR) and mid-infrared (MIR) frequency ranges, respectively. Residual magnetic fields affect the arc lamp, limiting reliable measurements below 40 $cm^{-1}$. Therefore, we primarily present results obtained with the Globar source. To detect the transmitted signal intensity in our measurements, we have used an IR Labs liquid-helium-cooled (and pumped) Si bolometer with cut-off filter for the range of 10 to 80  $cm^{-1}$, an IR Labs liquid-helium-cooled Si bolometer with cut-off filter for the range of 40 to 700  $cm^{-1}$, and an IR Labs liquid-helium-cooled Si bolometer with three different cut-off filters covering the range of 500 to 4000  $cm^{-1}$. The magnetic field of the cryostat can be tuned from -7 T to +7 T and was aligned parallel (perpendicular) to the light propagation vector (sample surface normal) in the Faraday configuration. The temperature range of the cryostat sample holder is from 4 K to about 300 K, while the sample is mounted with silver paste to the sample holder and situated in the vacuum chamber as the superconducting magnet during the measurement. The cryostat is positioned on an external linear translation stage, allowing the measurement of a sample in transmission (reflection) configuration and a bare hole (mirror) reference at any desired temperature, magnetic field, and incoming/outgoing polarization.
@@
Three types of experiments were done in this work: (i) Polarization-dependent transmission measurements using gold or KRS-5 (thallium bromoiodide) polarizers for the FIR or MIR frequency ranges, respectively, along the incoming and outgoing light beam paths. Our preliminary attempts to measure and fit non-polarized transmission data and, as a result, to extract the optical conductivity have failed, indicating the importance of the polarization dependence of the transmission spectrum and the anisotropy of the monoclinic crystal of  FePS3 in the magnetic state. Therefore, we have used an incoming polarizer for the transmission measurements. As a first stage, we measured the FIR and MIR spectra as a function of polarizer angle deep in the magnetic state, at 5 K, and at zero magnetic field (See App. A). We have identified two distinct transmission spectra separated by a $90^{\circ}$ polarizer angle difference, to which we refer here as "Pol 1" and "Pol 2". We emphasize that these two principal polarizations were not associated with the aba-b plane crystal axis using x-ray diffraction apparatus and are only determined by the optical transmission spectra. Following this criterion, we have used these two principal polarizations for all the temperature and magnetic field-dependent transmission measurements without an outgoing analyzer. (ii) Faraday rotation angle measurements with a rotating polarizer at the incoming optical path and a fixed analyzer at the outgoing optical beam. The fast protocol was used to investigate the temperature and magnetic field evolution of the Faraday angle rotation in FePS3. Since FePS3 has a monoclinic crystal structure and shows an anisotropic optical response in the antiferromagnetic state, the typical fast protocol procedure [22], taking into account only four measurements, i.e. combination of external magnetic field $\pm B$ and $A=P=\pm 45^{\circ}$, where $P$ and $A$ are the incoming polarizer and outgoing analyzer angles, respectively, was subject to possible birefringence artifacts. Therefore, we have conducted (iii) full protocol Faraday rotation angle measurements for each of the two principal polarizations used in this work. In the full protocol, the analyzer angle is fixed to one of the principal angles while measurement of the transmission intensity is being taken at intervals of $15^{\circ}$ for a full $360^{\circ}$ polarizer rotation. We have conducted the full protocol at a temperature of 5 K, i.e., deep in the antiferromagnetic state, and at fields of $B=\pm 7\leavevmode\nobreak\ T$.
II.3 Raman microscopy
Several samples of FePS3 were measured in Raman spectroscopy using a Labram HR Evolution Horiba system combined with a CryoVac liquid helium flow cryostat. A 532-nm unpolarized laser was used as the excitation source and was focused on the sample (inside the cryostat) through a 63× Olympus objective, and the laser power was set to 0.4 mW throughout the experiment. The scattering light from the sample was collected by the same objective, passed through the analyzer, and sent to a Czerny-Turner spectrometer equipped with either a 600 or 1800 grooves per mm grating and was detected by a liquid nitrogen-cooled CCD-array. The spectra were obtained as a function of temperature from 300 K down to 10 K at temperature intervals of 10 K, in addition to a measurement at 5 K, which is the base temperature of the Cryostat in our system. The 1800 grooves per mm grating was used to measure the Raman shift spectrum in the range of about $\pm 600$ $cm^{-1}$ of Raman shift (i.e. Stokes and anti-Stokes transitions) around the elastic line and to resolve the frequency resolution of the different excitations, in particular the modes around 80 to 130 $cm^{-1}$appearing below the phase transition of FePS3. The 600 grooves per mm grating was used to measure the large Raman shift range of up to approximately 4000 $cm^{-1}$in order to resolve the excitations seen in our data and in previous reports on FePS3 around 1000 $cm^{-1}$ (e.g, Ref. [23]).
II.4 Phonon calculation
In order to obtain the phonon frequencies of FePS3  we have performed first-principles Density-functional theory (DFT) calculations with the Quantum ESPRESSO PWscf (Plane-Wave Self-Consistent Field) package v6.6 [24, 25, 26]. Exchange–correlation effects were treated within the PBE (Perdew-Burke-Ernzerhof) functional and the generalized-gradient approximation (GGA) [27]. Scalar-relativistic pseudopotentials (PPs) consistent with PBE were used: PAW (Projector-Augmented Wave) for Fe and US (Ultrasoft) for P and S [28]. The conventional unit cell and lattice parameters were adopted from the parameters of FePS3 in the Materials Project database [29, 30]. To capture the antiferromagnetic ground state, we used collinear spin polarization ($n_{\mathrm{spin}}=2$) with two Fe sublattices initialized antiparallel (Fe1/Fe2 starting magnetization $\pm 0.3125$), and small seed magnetizations on P and S (0.1). The converged self-consistent solution has essentially zero net moment with total magnetization $\approx 0\leavevmode\nobreak\ \mu_{\mathrm{B}}$ per cell (absolute magnetization $\sim 15\leavevmode\nobreak\ \mu_{\mathrm{B}}$). Wavefunctions and charge density were expanded with kinetic-energy cutoffs of 90 Ry and 1080 Ry, respectively. Metallic occupations were treated using Marzari–Vanderbilt cold smearing with a width of 0.0147 Ry ($\approx$ 0.20 eV) [31]. Brillouin-zone sampling employed a Monkhorst–Pack $6\times 4\times 3$ mesh [32]. Symmetry was enabled and four symmetry operations (including inversion, two with fractional translations) were detected. Electronic self-consistency used plain mixing with a mixing parameter $\beta=0.4$ and a target energy change $\mathrm{conv_{thr}}=8\times 10^{-9}$ Ry; in this calculation, convergence was reached in 20 SCF iterations. Calculations were carried out for 20 atoms per cell with four atomic species. Phonons were computed within density-functional perturbation theory (DFPT) as implemented in QE’s ph.x [33] and resulted in zone-center ($\Gamma$-point) phonon frequencies.
@@ III.1 Infrared Transmission
In Figure 1, we show the transmission color map measured along the two principal axes. On one hand, the transmission spectrum at temperatures substantially higher than the Néel antiferromagnetic transition temperature, $T_{N}=118\leavevmode\nobreak\ K$, of  FePS3 shows no major difference between the two principal axes (See in App. A). On the other hand, a clear change in the spectra can be seen below the phase transition.
[…]
@@
Table 1: Comparison of theoretical (a our work, b adapted from Ref. [34]) and experimental (Pol 1 and Pol 2) vibrational modes in FePS3. Modes labeled with are identified as Raman active modes in theory (e.g. Ref. [34]) and observed in the experimental data. Modes labeled with are identified with an additional symmetry in Ref. [34]. Modes labeled with c were also observed experimentally in Ref. [35]. All values are in cm-1.
Symmetry Th.a Th.b Pol 1 Pol 2
$\mathrm{B}_{u}$ 75.8
$\mathrm{B}_{u}$ 101.9 102.1
$\mathrm{A}_{g}$$\mathrm{B}_{g}$ 107.7
$\mathrm{B}_{u}$ 119.4
$\mathrm{A}_{u}$ 129.7 128.5 133.1
$\mathrm{A}_{u}$ 139.0 140.8c 140.8c
@@
$\mathrm{A}_{u}$, $\mathrm{B}_{u}$ 153.8 150 153.1c
$\mathrm{A}_{g}$$\mathrm{B}_{g}$ 153
160.3c
$\mathrm{B}_{u}$ 161.7 158 163.5
$\mathrm{B}_{u}$ 167.6 165.2 165.9
$\mathrm{B}_{u}$ 174.5 175.7 174.5
$\mathrm{A}_{u}$ 178.0
$\mathrm{B}_{u}$ 179.3 184.7c 184.8c
$\mathrm{B}_{u}$ 202.5
$\mathrm{A}_{u}$ 202.7
$\mathrm{A}_{u}$ 207.5 207.4
$\mathrm{A}_{u}$, $\mathrm{B}_{u}$ 221.6 221 215.4c 215.1c
$\mathrm{B}_{u}$ 230.3 228.0 227.7
$\mathrm{B}_{u}$ 231.9 232.2
$\mathrm{B}_{u}$ 240.1
$\mathrm{A}_{u}$ 254.6 252
$\mathrm{B}_{u}$, $\mathrm{A}_{u}$ 254.6 258 257.4c
$\mathrm{B}_{u}$ 262 260.1
$\mathrm{B}_{u}$ 350.2 375.6
$\mathrm{B}_{u}$ 415.1 437
$\mathrm{B}_{u}$ 531.4 531.9
$\mathrm{A}_{u}$ 538.9
$\mathrm{B}_{u}$ 541.7 546.1
$\mathrm{B}_{u}$ 550.3
$\mathrm{A}_{u}$, $\mathrm{B}_{u}$ 556.7 570 556.3 555.0
Along with the drastic change of the overall transmission spectra below $T_{N}$, a new mode appears at about 122  $cm^{-1}$, or 15 meV, which seems to harden as the temperature is lowered. We will now focus on the spectral weight, temperature, magnetic field, and polarization-induced changes in the collective excitations spectrum, which are coupled to the magnetic order.